The doafter^{∗} package

Peter Schmitt^{†}  Mark Wooding

8 May 1996

Contents

1 Description 1 1.1 What it’s all about …. 1 1.2 Package options . . . . . . 2

2 Implementation 2 2.1 The main macro . . . . . 2 2.2 Test code . . . . . . . . . 5

A The GNU General Public Licence 7 A.1 Preamble . . . . . . . . . 8 A.2 Terms and conditions for copying, distribution and modification . . . . . . . . 8

A.3 Appendix: How to Apply These Terms to Your New Programs . . . . . . 12

1 Description

1.1 What it’s all about

\doafter It’s common for the TEX primitive \aftergroup to be used to ‘tidy up’ after a group. For example, L^{A}TEX’s colour handling uses this to insert appropriate \specials when the scope of a colour change ends. This causes several problems, though; for example, extra grouping must be added within boxes to ensure that the \specials don’t ‘leak’ out of their box and appear in odd places in the document. L^{A}TEX usually solves this problem by reading the box contents as an argument, although this isn’t particularly desirable. The \doafter macro provided here will solve the problem in a different way, by allowing a macro to regain control after all the \aftergroup things have been processed.

The macro works like this:

⟨doafter-cmd⟩ ::=

\doafter ⟨token⟩ ⟨group⟩

[figure]

The ⟨token⟩ can be any token you like, except an explicit braces, since it’s read as an undelimited macro argument. The ⟨group⟩ is a normal TEX group, surrounded by either implicit or explicit braces, or by \begingroup and \endgroup tokens.

∗The doafter package is currently at version 1.01, dated 8 May 1996.

†Peter came up with the basic implementation after I posed the problem in the comp.text.tex newsgroup. I fixed some really piddly little things, to improve it a bit, wrote the documentation, and turned the code into a nice docced package. Then Peter gave me an updated version, and I upgraded this from memory. Then he gave me some more tweaks which I haven’t incorporated.

Once the final closing token of the ⟨group⟩ is read, and any tokens saved up by \aftergroup have been processed, the ⟨token⟩ is inserted and processed. Under normal circumstances, this will be a macro.

There are some subtle problems with the current implementation, which you may need to be aware of:

• Since we’re inserting things after all the \aftergroup tokens, those tokens might read something they’re not expecting if they try to look ahead at the text after the group (e.g., with \futurelet). This is obviously totally unavoidable.

• Implicit braces (like \bgroup and \egroup) inserted using \aftergroup may be turned into explicit {_{1} and }_{2} characters within a \doafter group. This can cause probems under very specialised circumstances. The names \bgroup and \egroup are treated specially, and they will work normally (remaining as implicit braces). This should minimise problems caused by this slight difference. (This only applies to the last \aftergroup token in a group.)

• To handle the \aftergroup tokens properly, \doafter has to insert some \aftergroup tokens of its own. It will then process the other tokens some more, and set them up to be read again. This does mean that after the group ends, some assignments and other ‘stomach operations’ will be performed, which may cause problems in alignments and similar places.

1.2 Package options

There are a fair few docstrip options provided by this packge:

driver extracts the documentation driver. This isn’t usually necessary.

package extracts the code as a standalone package, formatted for either L^{A}TEX 2ε or Plain TEX.

latex2e inserts extra identification code for a L^{A}TEX 2εpackage.

plain inserts some extra code for a Plain TEX package.

macro just extracts the raw code, for inclusion in another package.

test extracts some code for testing the current implementation.

2 Implementation

2.1 The main macro

We start outputting code here. If this is a Plain TEX package, we must make ‘@’ into a letter.

1 ⟨∗macro | package⟩

2 ⟨+plain⟩\catcode`\@=11

\doafter The idea is to say \doafter ⟨token⟩ ⟨group⟩ and expect the ⟨token⟩ to be processed after the group has finished its stuff, even if it contains \aftergroup things. My eternal gratitude goes to Peter Schmitt, who came up with most of the solution implemented here; I’ve just tidied up some very minor niggles and things later.

Let’s start with some preamble. I’ll save the (hopefully) primitive \aftergroup in a different token.

 3 \let\@@aftergroup\aftergroup

Now to define the ‘user’ interface. It takes a normal undelimited argument, although this must be a single token; otherwise eveything will go wrong. It assumes that the token following is some kind of group opening thing (an explicit or implicit character with catcode 1, or a \begingroup token). To make this work, I’ll save the token, together with an \@@aftergroup (to save an \expandafter later) in a temporary macro which no-one will mind me using, and then look ahead at the beginning-group token.

 4 \def\doafter#1{%
 5 \def\@tempa{\@@aftergroup#1}%
 6 \afterassignment\doafter@i\let\@let@token%
 7 }

I now have the token in \@let@token, so I’ll put that in. I’ll then make \aftergroup do my thing rather than the normal thing, and queue the tokens \@prepare@after and the \doafter argument for later use.

 8 \def\doafter@i{%
 9 \@let@token%
10 \let\aftergroup\@my@aftergroup%
11 \@@aftergroup\@prepare@after\@tempa%

12 }

\@my@aftergroup Now the cleverness begins. We keep two macros (Peter’s original used count registers) which keep counts of the numbers of \aftergroups, both locally and globally. Let’s call the local counter n and the global one N. Every time we get a call to our \aftergroup hack, we set n := n + 1 and N := n, and leave the token given to us for later processing. When we actually process an \aftergroup token properly, set N := N − 1 to indicate that it’s been handled; when they’re all done, we’ll have N = n, which is exactly what we’d have if there weren’t any to begin with.

13 \def\ag@cnt@local{0 }
14 \let\ag@cnt@global\ag@cnt@local

Now we come to the definition of my version of \aftergroup. I’ll just add the token \@after@token before every \aftergroup token I find. This means there’s two calls to \aftergroup for every one the user makes, but these things aren’t all that common, so it’s OK really. I’ll also bump the local counter, and synchronise them.

15 \def\@my@aftergroup{%
16 \begingroup%
17  \count@\ag@cnt@local%
18  \advance\count@\@ne%
19  \xdef\ag@cnt@global{\the\count@\space}%
20 \endgroup%

21 \let\ag@cnt@local\ag@cnt@global%
22 \@@aftergroup\@after@token\@@aftergroup%

23 }

Now what does \@after@token we inserted above actually do? Well, this is more exciting. There are actually two different variants of the macro, which are used at different times.

\@after@token The default \@after@token starts a group, which will ‘catch’ \aftergroup tokens which I throw at it. I put the two counters into some scratch count registers. (There’s a slight problem here: Plain TEX only gives us one. For the sake of evilness I’ll use \clubpenalty as the other one. Eeeek.) I then redefine \@after@token to the second variant, and execute it. The \@start@after@group macro starts the group, because this code is shared with \@prepare@after below.

24 \def\@after@token{%
25 \@start@after@group%
26 \@after@token%

27 }

28 \def\@start@after@group{%
29 \begingroup%
30 \count@\ag@cnt@global%
31 \clubpenalty\ag@cnt@local%
32 \let\@after@token\@after@token@i%

33 }

\@after@token@i I have \count@ = N and \@tempcnta = n. I’ll decrement N, and if I have N = n, I know that this is the last token to do, so I must insert an \@after@all after the token. This will close the group, and maybe insert the original \doafter token if appropriate.

34 \def\@after@token@i{%
35 \advance\count@\m@ne%
36 \ifnum\count@=\clubpenalty%
37  \global\let\ag@cnt@global\ag@cnt@local%
38  \expandafter\@after@aftertoken\expandafter\@after@all%
39 \else%
40  \expandafter\@@aftergroup%
41 \fi%

42 }

Finally, establish a default meaning for \@after@all.

43 \let\@after@all\endgroup

\@prepare@after If this group is handled by \doafter, then the first \aftergroup token isn’t

\@after@token; it’s \@prepare@after.

There are some extra cases to deal with:

• If N = n then there were no \aftergroup tokens, so we have an easy job. I’ll just let the token do its stuff directly.

• Otherwise, N > n, and there are \aftergroup tokens. I’ll open the group, and let \@after@token do all the handling.

44 \def\@prepare@after{%
45 \ifx\ag@cnt@local\ag@cnt@global\else%
46  \expandafter\@prepare@after@i%
47 \fi%

48 }

49 \def\@prepare@after@i#1{%
50 \@start@after@group%
51 \def\@after@all{\@@aftergroup#1\endgroup}%

52 }

\@after@aftertoken This is where all the difficulty lies. The next token in the stream is an \aftergroup one, which could be more or less anything. We have an argument, which is some code to do after the token has been \aftergrouped.

If the token is anything other than a brace (i.e., an explicit character of category 1 or 2) then I have no problem; I can scoop up the token with an undelimited macro argument. But the only way I can decide if this token is a brace (nondestructively) is with \futurelet, which makes the token implicit, so I can’t decide whether it’s really dangerous.

There is a possible way of doing this^{1} which relates to nobbling the offending token with \string and sifting through the results. The problem here involves scooping up all the tokens of a \stringed control sequence, which may turn out to be ‘\csname\endcsname’ or something equally horrid.

The solution I’ve used is much simpler: I’ll change \bgroup and \egroup to stop them from being implicit braces before comparing.

53 \def\@after@aftertoken#1{%
54 \let\bgroup\relax\let\egroup\relax%
55 \toks@{#1}%
56 \futurelet\@let@token\@after@aftertoken@i%

57 }

58 \def\@after@aftertoken@i{%
59 \ifcat\noexpand\@let@token{%
60  \@@aftergroup{%
61 \else\ifcat\noexpand\@let@token}%
62  \@@aftergroup}%
63 \else%
64  \def\@tempa##1{\@@aftergroup##1\the\toks@}%
65  \expandafter\expandafter\expandafter\@tempa%
66 \fi\fi%

67 }

Phew!

68 ⟨+plain⟩\catcode`\@=12

69 ⟨/macro | package⟩

2.2 Test code

The following code gives \doafter a bit of a testing. It’s based on the test suite I gave to comp.text.tex, although it’s been improved a little since then.

1Due to Peter Schmitt, again.

The first thing to do is define a control sequence with an ‘@’ sign in its name, so we can test catcode changes. This also hides an \aftergroup within a macro, making life more difficult for prospective implementations.

70 ⟨∗test⟩

71 \catcode`\@=11
72 \def\at@name{\aftergroup\saynine}
73 \def\saynine{\say{ix}}
74 \catcode`\@=12

Now define a command to write a string to the terminal. The name will probably be familiar to REXX hackers.

75 \def\say{\immediate\write16}

Test one: This is really easy; it just tests that the thing works at all. If your implementation fails this, it’s time for a major rethink.

76 \say{Test one... (1--2)}
77 \def\saytwo{\say{ii}}
78 \doafter\saytwo{\say{i}}

Test two: Does \aftergroup work?

79 \say{Test two... (1--4)}
80 \def\saythree{\say{iii}}
81 \def\sayfour{\say{iv}}
82 \doafter\sayfour{\say{i}\aftergroup\saythree\say{ii}}

Test three: Test braces and \iffalse working as they should. Several proposed solutions based on \writeing the group to a file get upset by this test, although I forgot to include it in the torture test. It also tests whether literal braces can be \aftergrouped properly. (Added a new test here, making sure that \bgroup is left as an implicit token.)

83 \say{Test three... (1--4, `\string\bgroup', 5)}
84 \def\sayfive{\say{v}}
85 \doafter\sayfive{%
86 \say{i}%
87 \aftergroup\say%
88 \aftergroup{%
89 \aftergroup\romannumeral\aftergroup3%
90 \aftergroup}%
91 \iffalse}\fi%
92 \aftergroup\def%
93 \aftergroup\sayfouretc%
94 \aftergroup{%
95 \aftergroup\say%
96 \aftergroup{%
97 \aftergroup i%
98 \aftergroup v%
99 \aftergroup}%
100 \aftergroup\say%
101 \aftergroup{%
102 \aftergroup\string%
103 \aftergroup\bgroup%
104 \aftergroup}%
105 \aftergroup}%

106 \aftergroup\sayfouretc%
107 \say{ii}%

108 }

Test four: Make sure the implementation isn’t leaking things. This just makes sure that \aftergroup is its normal reasonable self.

109 \say{Test four... (1--3)}
110 {\say{i}\aftergroup\saythree\say{ii}}

Test five: Nesting, aftergroup, catcodes, grouping. This is the ‘torture’ test I gave to comp.text.tex, slightly corrected (oops) and amended. It ensures that nested groups and \doafters work properly (the latter is actually more likely than might be imagined).

111 \say{Test five... (1--14)}
112 \def\sayten{\say{x}}
113 \def\saythirteen{\say{xiii}}
114 \def\sayfourteen{\say{xiv}}
115 \doafter\sayfourteen\begingroup%
116 \say{i}%
117 {\say{ii}\aftergroup\sayfour\say{iii}}%
118 \def\saynum{\say{viii}}%
119 \doafter\sayten{%
120  \say{v}%
121  \def\saynum{\say{vii}}%
122  \catcode`\@=11%
123  \aftergroup\saynum%
124  \say{vi}%
125  \at@name%
126  \saynum%

127 }%

128 \say{xi}%
129 \aftergroup\saythirteen%
130 \say{xii}%
131 \endgroup
132 \end

133 ⟨/test⟩

That’s it. All present and correct.

Appendix

A The GNU General Public Licence

The following is the text of the GNU General Public Licence, under the terms of which this software is distrubuted.

GNU GENERAL PUBLIC LICENSE Version 2, June 1991

Copyright (C) 1989, 1991 Free Software Foundation, Inc.

675 Mass Ave, Cambridge, MA 02139, USA Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

A.1 Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

A.2 Terms and conditions for copying, distribution and modification

0. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”.

Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

1. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program.

You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions:

(a) You must cause the modified files to carry prominent notices stating that you changed the files and the date of any change.

(b) You must cause any work that you distribute or publish, that in whole or in part contains or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

(c) If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it.

Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program.

In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following:

(a) Accompany it with the complete corresponding machine-readable source code, which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

(b) Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

(c) Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable.

If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

4. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

5. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

6. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

7. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program.

If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances.

It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice.

This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

8. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

9. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

10. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

11. Because the Program is licensed free of charge, there is no warranty for the Program, to the extent permitted by applicable law. except when otherwise stated in writing the copyright holders and/or other parties provide the program “as is” without warranty of any kind, either expressed or implied, including, but not limited to, the implied warranties of merchantability and fitness for a particular purpose. The entire risk as to the quality and performance of the Program is with you. Should the Program prove defective, you assume the cost of all necessary servicing, repair or correction.

12. In no event unless required by applicable law or agreed to in writing will any copyright holder, or any other party who may modify and/or redistribute the program as permitted above, be liable to you for damages, including any general, special, incidental or consequential damages arising out of the use or inability to use the program (including but not limited to loss of data or data being rendered inaccurate or losses sustained by you or third parties or a failure of the Program to operate with any other programs), even if such holder or other party has been advised of the possibility of such damages.

END OF TERMS AND CONDITIONS

A.3 Appendix: How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

<one line to give the program's name and a brief idea of what it does.>
Copyright (C) 19yy  <name of author>
This program is free software; you can redistribute it and/or modify
it under the terms of the GNU General Public License as published by
the Free Software Foundation; either version 2 of the License, or
(at your option) any later version.
This program is distributed in the hope that it will be useful,
but WITHOUT ANY WARRANTY; without even the implied warranty of
MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
GNU General Public License for more details.
You should have received a copy of the GNU General Public License
along with this program; if not, write to the Free Software
Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.

Also add information on how to contact you by electronic and paper mail. If the program is interactive, make it output a short notice like this when it

starts in an interactive mode:

Gnomovision version 69, Copyright (C) 19yy name of author
Gnomovision comes with ABSOLUTELY NO WARRANTY; for details type `show w'.
This is free software, and you are welcome to redistribute it
under certain conditions; type `show c' for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items–whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright interest in the program
`Gnomovision' (which makes passes at compilers) written by James Hacker.
<signature of Ty Coon>, 1 April 1989
Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

Index

The italic numbers denote the pages where the corresponding entry is described, numbers underlined point to the definition, all others indicate the places where it is used.

Symbols

\@ . . . . . . . . . . . . . . . 2, 68, 71, 74, 122 \@@aftergroup . . . . . . . . . . . . . . . .

… 3, 5, 11, 22, 40, 51, 60, 62, 64

\@after@aftertoken . . . . . . . . . . 38, \underline{53}
\@after@aftertoken@i . . . . . . . . 56, 58

\@after@all . . . . . . . . . . . . . 38, 43, 51 \@after@token . . . . . . . . . . . . . . 22, \underline{24} \@after@token@i . . . . . . . . . . . . 32, \underline{34} \@let@token . . . . . . . . . 6, 9, 56, 59, 61 \@my@aftergroup . . . . . . . . . . . . 10, \underline{13} \@prepare@after . . . . . . . . . . . . 11, \underline{44}

\@prepare@after@i . . . . . . . . . . 46, 49
\@start@after@group . . . . . . 25, 28, 50

\@tempa . . . . . . . . . . . . . . 5, 11, 64, 65

A

\afterassignment . . . . . . . . . . . . . . 6

\aftergroup . . . . . . . . 3, 10, 72, 82,

87–90, 92–106, 110, 117, 123, 129 \ag@cnt@global . . 14, 19, 21, 30, 37, 45 \ag@cnt@local 13, 14, 17, 21, 31, 37, 45 \at@name . . . . . . . . . . . . . . . . . 72, 125

C

\clubpenalty . . . . . . . . . . . . . . . 31, 36 \count@ . . . . . . . . . . . 17–19, 30, 35, 36

D

\doafter …. 1, \underline{3}, 78, 82, 85, 115, 119 \doafter@i . . . . . . . . . . . . . . . . . 6, 8

E \end . . . . . . . . . . . . . . . . . . . . . . . . 132

F \futurelet . . . . . . . . . . . . . . . . . . 56

R \romannumeral . . . . . . . . . . . . . . . . 89

S \say . . . . . . . . . . . . . . . . 73, 75–84,

86, 87, 95, 100, 107, 109–114, 116–118, 120, 121, 124, 128, 130

\sayfive . . . . . . . . . . . . . . . . . . 84, 85 \sayfour . . . . . . . . . . . . . . 81, 82, 117 \sayfouretc . . . . . . . . . . . . . . . 93, 106 \sayfourteen . . . . . . . . . . . . . 114, 115 \saynine . . . . . . . . . . . . . . . . . . 72, 73 \saynum . . . . . . . . . . 118, 121, 123, 126 \sayten . . . . . . . . . . . . . . . . . 112, 119 \saythirteen . . . . . . . . . . . . . 113, 129 \saythree . . . . . . . . . . . . . 80, 82, 110 \saytwo . . . . . . . . . . . . . . . . . . . 77, 78

T \toks@ . . . . . . . . . . . . . . . . . . . . 55, 64